OpenStudy (anonymous):

5-2x3+6(3(squared)+1)

OpenStudy (jhannybean):

\(5-(2\cdot 3)+6(3^2+1)\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\ 5-(2\cdot 3)+6(3^2+1) = 5-6+6(9+1)=-1+6(10)=-1+60=59\]